A rowing team rowing with the current traveled 20 kn in 2 hrs. Rowing against the current, the team rowed 12 kn in the same amount of time. Find the rate of the team in calm water and the rate of the current.
Answer by scott8148(6628)

let x="rate of team" and y="rate of current" __ d=rt

"A rowing team rowing with the current traveled 20 kn in 2 hrs" __ 20=2(x+y) __ 10=x+y

"Rowing against the current, the team rowed 12 kn in the same amount of time" __ 12=2(x-y) __ 6=x-y

adding equations __ 10+6=x+y+x-y __ 16=2x __ 8=x

substituting __ 10=8+y __ 2=y